OpenStudy (anonymous):

what letter grade is a 15 out of 22

OpenStudy (zale101):

15/22=0.68181818181818181818181818181818 0.68

OpenStudy (anonymous):

what does this mean?

OpenStudy (anonymous):

what letter grade is a 15/22? a C, D?

OpenStudy (anonymous):

maybe a c- or d .

OpenStudy (anonymous):

100- 90 = A 89- 80 = B 79- 70 = C 69- 60 = D 59 - below = F

OpenStudy (zale101):

it depends on your school's grading system, every teacher has his/her unique grading style

OpenStudy (anonymous):

ok thanks! ;)

OpenStudy (anonymous):

A 68 is a D or D+ depends on how your teacher grades :)

OpenStudy (zale101):

No Problem :>

OpenStudy (anonymous):

Also, what is mechanical advantage? I don't get that...

OpenStudy (zale101):

idk :/

OpenStudy (anonymous):

It is used to describe an amount of force